OpenStudy (anonymous):

solve 4/5 x - 3 3/4x = 2 1/2

OpenStudy (anonymous):

.8x - 3.75x = 2.5 -2.95x = 2.5 x = -.847

OpenStudy (anonymous):

\[\frac{4 x}{5}-\frac{15 x}{4}\text{ = }\frac{5}{2}\]\[-\frac{59 x}{20}\text{ = }\frac{5}{2} \]\[x=-\frac{50}{59} \] x = -0.8474576271186440677966101694915254237288135593220338983050

OpenStudy (anonymous):

x has 58 non repeating digits.